OpenStudy (anonymous):

A die is rolled three times. Find the chance that three different faces appear

OpenStudy (anonymous):

Hi in three throws : 6*6*6=216 three different faces : \[\left(\begin{matrix}6 \\1\end{matrix}\right) *\left(\begin{matrix}5 \\1\end{matrix}\right)*\left(\begin{matrix}4 \\ 1\end{matrix}\right)\] =120 -->p=120/216

OpenStudy (anonymous):

thanks